OpenStudy (soahelp):

MEDAL!!!!!!!!! Meg is 6 years older than Victor. Meg's age is 2 years less than five times Victor's age. The equations below model the relationship between Meg's age (m) and Victor's age (v): m = v + 6 m = 5v − 2 Which is a possible correct method to find Meg's and Victor's ages?

OpenStudy (soahelp):

Solve m + 6 = 5m − 2 to find the value of m. Write the points where the graphs of the equations intersect the x axis. Solve v + 6 = 5v − 2 to find the value of v. Write the points where the graphs of the equations intersect the y axis.

OpenStudy (anonymous):

You would set the two equations equal to each other so v+6=5v-2 and solve for v then plug that back into the first equation to solve for m, v and m are the ages